# First Measurement of the Decay Asymmetry in the pure W-boson-exchange Decay $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$

### 18 Jan 2024 v3

@@ Abstract
Körner (1992), CCQM [7] 2.6 0 - - -
Xu (1992), Pole [8] 1.0 0 00 $7.94$
Źencaykowski (1994), Pole [9] 3.6 0 - - -
Ivanov (1998), CCQM [10] 3.1 0 - - -
Sharma (1999), CA [11] 1.3 0 - - -
Geng (2019), SU(3) [12] $5.7\pm 0.9$ $0.94^{+0.06}_{-0.11}$ $2.7\pm 0.6$ $16.1\pm 2.6$
Zou (2020), CA [6] $7.1$ $0.90$ $4.48$ $12.10$
Zhong (2022), SU(3)${}^{a}$ [13] $3.8^{+0.4}_{-0.5}$ $0.91^{+0.03}_{-0.04}$ $3.2\pm 0.2$ $8.7^{+0.6}_{-0.8}$
Zhong (2022), SU(3)${}^{b}$ [13] $5.0^{+0.6}_{-0.9}$ $0.99\pm 0.01$ $3.3^{+0.5}_{-0.7}$ $12.3^{+1.2}_{-1.8}$
BESIII (2018) [14] $5.90\pm 0.86\pm 0.39$
PDG Fit (2022) [2] $5.5\pm 0.7$

### 20 Jan 2024 v4

@@ Abstract
Körner (1992), CCQM [7] 2.6 0 $\cdots$ $\cdots$ $\cdots$
Xu (1992), Pole [8] 1.0 0 00 $7.94$cdots
Źencaykowski (1994), Pole [9] 3.6 0 $\cdots$ $\cdots$ $\cdots$
Ivanov (1998), CCQM [10] 3.1 0 $\cdots$ $\cdots$ $\cdots$
Sharma (1999), CA [11] 1.3 0 $\cdots$ $\cdots$ $\cdots$
Geng (2019), SU(3) [12] $5.7\pm 0.9$ $0.94^{+0.06}_{-0.11}$ $2.7\pm 0.6$ $16.1\pm 2.6$cdots
Zou (2020), CA [6] $7.1$ $0.90$ $4.48$ $12.10$cdots
Zhong (2022), SU(3)${}^{a}$ [13] $3.8^{+0.4}_{-0.5}$ $0.91^{+0.03}_{-0.04}$ $3.2\pm 0.2$ $8.7^{+0.6}_{-0.8}$cdots
Zhong (2022), SU(3)${}^{b}$ [13] $5.0^{+0.6}_{-0.9}$ $0.99\pm 0.01$ $3.3^{+0.5}_{-0.7}$ $12.3^{+1.2}_{-1.8}$cdots
BESIII (2018) [14] $5.90\pm 0.86\pm 0.39$cdots $\cdots$ $\cdots$ $\cdots$
PDG Fit (2022) [2] $5.5\pm 0.7$cdots $\cdots$ $\cdots$ $\cdots$